# No more hard prompts: SoftSRV prompting for synthetic data generation

### v2

@@ 3 Empirical Evaluation
In the our empirical evaluation of the SoftSRV framework, we consider a supervised fine-tuning setting where a small Gemma 2B model (Team et al., 2024) is fine-tuned on synthetic data generated by a larger decoder-only language model across several different benchmark datasets.
@@ 3.1 Domains and datasets
In order to demonstrate the generality of the proposed approach, we consider fine-tuning for several disparate domains (coding, mathematics, reasoning) using the same exact pipeline with no particular specialization to any of the particular domains. We briefly describe the specific benchmark we use from each domain.
@@
Reasoning – BoolQ (Clark et al., 2019). For the general reasoning domain, we consider the binary question answering dataset from the BoolQ benchmark. The questions arise organically from anonymized Google search queries which can be answered as either ‘true’ or ‘false’. Each question and answer is paired with a passage (average length of 108 tokens) that is extracted from a relevant Wikipedia page. This is evaluated as a scoring/classification task and accuracy is reported.
@@
The above chosen benchmarks aim to cover a wide variety of tasks, each with varying degrees of complexity – both in terms of solving the task and in terms of generating synthetic data for the task. The MBBP task requires basic Python programming knowledge to solve, but the questions are generally short, follow a similar pattern and are concerned with a relatively narrow set of themes. The GSM8K task requires basic math and language comprehension skills, but generating the problem is arguably even harder than solving it. It requires generating a premise containing several numerical quantities and then a question that can be answered using the provided information in a non-trivial fashion. Nonetheless the premises are somewhat formulaic and thematically similar. The BoolQ reasoning task, which requires general reading comprehension to solve, is perhaps the most difficult task to generate synthetic data for. Generating a problem requires writing a long (relative to MBPP and GSM8K) passage on an arbitrary topic that contains a collection of facts, but that does not necessarily stick to any formula or theme, and then generate a true/false question that can be answered directly by the passage. As we shall see in the empirical evaluation that follows, the level of difficulty in generating a high-quality question can impact the relative quality and value of generated synthetic data.
3.2 Hard-prompting baselines
Typical hard prompt engineering approaches involve manually creating prompt templates that are then seeded with text from the desired target domain. To give a (simplistic) illustrative example, a template could be:
Consider the following [article], write a textbook quality summary of the topic suitable for a high-school audience,
where the placeholder [article] would be replaced with example texts from the desired domain, producing several distinct prompts. In this study, we consider the following two hard-prompting variants.
The first, denoted simply as hard prompt ($\mathrm{HP}$), uses a domain-specific hard prompt template to generate a question followed by another domain specific hard prompt template to generate answers (the detailed workflow is discussed in Subsection 3.3). In Appendix B, we provide the exact templates used by the $\mathrm{HP}$ method. To reach these templates, we undertook several iterations of hard prompt engineering and reported the result of the best performing method. In particular, we found that prompting for a "diverse" set of questions was crucial (a comparison plot is presented in Appendix A.5).
The second approach, hard-prompting with self-refinement ($\mathrm{HP}_{\mathrm{SR}}$), similarly uses a hard prompt template to generate questions but also iteratively conducts several rounds of self-critique to improve or accept the question (Madaan et al., 2023). Again, critique and refinement prompts are in Appendix B.
3.3 Empirical evaluation procedure
For each benchmark dataset and each data generation approach, the evaluation pipeline is as follows.
1. Train SoftSRV prompts. For the SoftSRV methods, we first train the soft prompt parameters with a frozen large decoder-only LM backbone using the questions found in the training fold of the dataset, which serves as our sample from the target distribution. Recalling we embed each question, $\mathbf{z}_{i}=\mathrm{emb}(x_{i})$, we run an Adam optimizer to minimize the causal next-word-prediction loss, $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(\mathbf{z}_{i}),x_{i})$, where $\mathbf{P}_{\theta}(\mathbf{z})$ is the conditioned soft prompt and $H$ is the frozen LLM (post input embedding layer). The sequence embeddings, $\mathrm{emb}(\cdot)$, are computed as the average of token embeddings computed by a small off-the-shelf decoder-only LM. This simple embedding approach is used to both limit the amount of additional computation, but also to ensure the embedding is somewhat lossy in order to make the reconstruction task (i.e. minimizing $\ell(H(\mathbf{P}_{\theta}(\mathrm{emb}(x_{i})),x_{i}$) challenging. The simpler $\mathrm{SS}_{\mathrm{NSP}}$ method does not use this sequence embedding as it operates with a non-contextual soft prompt.
Since we seek an automated hands off approach, we avoid any domain-specific hyperparameter tuning. Specifically, for all SoftSRV variants and all benchmarks, the length of the prompt $t$ is fixed to be 128, the number of training steps was set to 20K, and the learning rate is fixed to $5\mathrm{e}{-6}$, which we found to be reasonable defaults. The $\mathrm{SS}_{\mathrm{MC}}$ method uses MLPs with 3 feed forward layers and 128 hidden dimensions. For the $\mathrm{SS}_{\mathrm{MPk}}$ variant, we primarily evaluate with $k=2$ to limit to the total number of parameters, although a partial exploration for other values of $k$ is presented in Appendix A.4.
@@ 3.3 Empirical evaluation procedure
2. Generate Questions. Once trained, we generate synthetic questions with the SoftSRV model. That is, we pass in the questions from the training dataset, $x_{i}$, and produce a new sequence $x_{i}^{\prime}\sim H(\mathbf{P}_{\theta}(\mathrm{emb}(x_{i})))$ via temperature sampling (with default temp=1). For the $\mathrm{SS}_{\mathrm{NSP}}$ variant, no context vector and, thus, no training examples are used during generation. For all SoftSRV methods, no hard prompt template of any kind is used.

### v3

@@ 3 Empirical Evaluation
To empirically evaluate the SoftSRV framework, we consider a supervised fine-tuning setting where we fine-tune a small Gemma 2B model (Team et al., 2024) using synthetic data generated by a large foundational decoder-only language model. To generate the data, we only consider methods that directly prompt the larger model and/or train a relatively small number of parameters (as in the case of SoftSRV). We do not assume the ability to conduct full fine-tuning of the large foundational model, as it is generally computationally prohibitive.
@@ 3.1 Domains and Datasets
To demonstrate the generality of the proposed approach, we generate fine-tuning data for several disparate domains (coding, mathematics, reasoning) using the same exact pipeline with no particular specialization to any of the domains.
Code – MBPP (Austin et al., 2021) consists of short Python programming exercises (metric: 3-shot pass@1).
Math – GSM8K (Cobbe et al., 2021) contains grade school math word problems (metric: 5-shot pass@1).
Reasoning – BoolQ (Clark et al., 2019) is a question answering dataset where each true/false question is paired with a passage (metric: accuracy).
The above benchmarks aim to cover a wide variety of tasks, each with varying degrees of complexity – both in terms of solving the task and in terms of generating synthetic data. Please see Appendix A.1 for a detailed discussion.
3.2 Baselines
Typical prompt engineering approaches for synthetic text generation involve manually creating natural language prompt templates that are then seeded with text from the desired target domain, usually taken from the training set. To give a simple illustrative example, a template could be:
Consider the following [article], write a summary of the topic suitable for a high-school audience,
where the placeholder [article] would be replaced with example texts from training fold, producing several distinct prompts. In this study, we consider the following two natural language prompt template variants.
The first, denoted simply as prompt template ($\mathrm{PT}$), uses a domain-specific template to generate a question followed by another domain specific template to generate answers (the detailed workflow is discussed in Subsection 3.3). In Appendix C, we provide the exact templates used by the $\mathrm{PT}$ method. To design these templates, we undertook several iterations of prompt engineering and reported the result of the best performing method. In particular, we found that prompting for a "diverse" set of questions was crucial (a comparison plot is presented in Appendix A.4).
The second approach, prompt template with self-refinement ($\mathrm{PT}_{\mathrm{SR}}$), similarly uses a template to generate questions but also iteratively conducts several rounds of self-critique to improve or accept the question (Madaan et al., 2023). Again, critique and refinement prompts are shown in Appendix C.
We focus on comparing to natural language template prompting approaches as they are currently the most effective, general, and widely used approaches for practical synthetic data generation (see related work in Section 4).
3.3 Empirical Evaluation Procedure
Below, we describe the data generation and evaluation pipeline used for each benchmark dataset and each data generation approach. Figure 2 also provides a diagram, illustrating the data generation pipeline step used by SoftSRV and the prompt template baseline approaches.
Figure 2: An illustration of the workflow for generating synthetic SFT data with SoftSRV (top panel) and natural language prompt templates (bottom panel); step numbering matches the discussion in Section 3.3. For SoftSRV, we use questions from the train set to (1) train a contextual embedding, (2) embed the same training sequences to serve as context vector to the trained embedding which is then fed to the LLM to generate synthetic questions, and (3) we generate answers by simply feeding the synthetic questions to an LLM. In the baseline prompt template framework, we have no training step (1) albeit there is offline human effort needed to generate the various prompt templates for each data domain; step (2) generates questions, using questions from the train set to fill the natural language template (optionally conduct rounds of refinement prompting), and (3) generates answers to these questions again using a template, but populated with the synthetic questions.
Step 1. Train Parameters (SoftSRV only). For the SoftSRV methods, we first train its parameters with a frozen large decoder-only LM backbone using the questions found in the training fold of the benchmark dataset, which serves as our sample from the target distribution. We embed each question, $\mathbf{z}_{i}=E(x_{i})$, and run an Adam optimizer to minimize the causal next-word-prediction loss, $\mathrm{argmin}_{\theta}\sum_{i=1}^{n}\ell(H(\mathbf{P}_{\theta}(\mathbf{z}_{i}),x_{i})$, where $\mathbf{P}_{\theta}(\mathbf{z})$ is the contextual SoftSRV embedding and $H$ is the frozen LLM (post input embedding layer). The sequence embeddings, $E(\cdot)$, are set to be the average of token embeddings computed by a small off-the-shelf decoder-only LM. This simple choice for $E(\cdot)$ is used to both limit the amount of additional computation, but also to ensure the sequence embedding is somewhat lossy in order to make the reconstruction task (i.e. minimizing $\ell(H(\mathbf{P}_{\theta}(E(x_{i})),x_{i}$) challenging). The simpler $\mathrm{SS}_{\mathrm{NP}}$ method does not use this sequence embedding as it is not a contextual parameterization.
Since we are interested in a low-touch general framework, we avoid any domain-specific hyperparameter tuning for SoftSRV. Specifically, for all SoftSRV variants and all benchmarks, the length of the prompt $t$ is fixed to be 128, the number of training steps was set to 20K, and the learning rate is fixed to $5\mathrm{e}{-6}$, which we found to be reasonable defaults. The $\mathrm{SS}_{\mathrm{MC}}$ method uses MLPs with 3 feed forward layers and 128 hidden dimensions. For the $\mathrm{SS}_{\mathrm{MPk}}$ variant, we primarily evaluate with $k=2$ to limit to the total number of parameters, although a partial exploration for other values of $k$ is presented in Appendix A.6.
Step 2. Generate Questions. The main application of SoftSRV is in this step, where we generate new questions for the target fine-tune task. Both the SoftSRV and the prompt template methods use all examples in the train data during this question generation phase. Specifically, SoftSRV uses examples $x_{i}$ from the train fold as context to generate a synthetic questions $x_{i}^{\prime}\sim H(\mathbf{P}_{\theta}(E(x_{i})))$, while $\mathrm{PT}$ and $\mathrm{PT}_{\mathrm{SR}}$ use $x_{i}$ to populate a template that is then passed to the LLM. Both approaches leverage temperature sampling. We generate 50,000 questions for MBPP and GSM8K and 20,000 questions for BoolQ (see Appendix D for examples). Appendix A.2 contains additional details on this step.
Step 3. Generate Answers. After generating the questions, all methods essentially follow the same procedure to generate answers using an off-the-self LLM. The only difference being, in the case of SoftSRV, we pass the question directly to the LLM without any domain specific prompting to preserve its domain agnostic nature. In the case of the prompt template baselines, we use a domain specific template combined with the generated question to query the off-the-shelf LLM for an answer. For all methods, once we have full (questions, answer) fine-tuning examples generated, we run a decontamination process to remove any examples that may have been inadvertently leaked to the pretrained LLM, as is standard practice (details provided in Appendix A.3).
Step 4. Fine-tune and Evaluate Downstream Model. Finally, for all methods, we use the generated (question, answer) pairs to fine-tune the target 2B Gemma 2 model. We use a batch-size of 16 with sequence length 8192 and with a learning rate with linear warmup from 0 to 1e-6 over 100 steps, followed by a cosine annealing schedule. We evaluate the performance of these fine-tuned models on the test fold of the respective benchmark using the metrics in Section 3.1.
3.4 Comparing our Parameterizations of SoftSRV
Table 1: Downstream task performance of Gemma 2B models fine-tuned on the data generated by each of the three SoftSRV methods, evaluated at the 2K checkpoint. The number of non-synthetic examples used as seed data for the prompt templates and SoftSRV models is reported as $N_{r}$.
task $N_{r}$ $\mathrm{SS}_{\mathrm{NP}}$ $\mathrm{SS}_{\mathrm{MP2}}$ $\mathrm{SS}_{\mathrm{MC}}$
MBPP 384 0.296 0.312 0.372
GSM8K 7,473 0.413 0.401 0.463
BoolQ 9,427 0.806 0.825 0.831
We start our empirical analysis by comparing different SoftSRV parameterizations to identify the highest performing configuration. Here, two salient questions are 1) Which parameterization most is effective and 2) whether it is essential to have a contextual SoftSRV embedding that leverages the context vector as opposed to a non-contextual SoftSRV embedding. In Table  1, we present the performance of Gemma 2B fine-tuned on synthetic datasets generated by $\mathrm{SS}_{\mathrm{NP}}$, $\mathrm{SS}_{\mathrm{MP2}}$ and $\mathrm{SS}_{\mathrm{MC}}$, respectively. These results show that the non-contextual method, $\mathrm{SS}_{\mathrm{NP}}$, generates synthetic fine-tuning data that results in lower performance than the other (contextual) parameterizations, indicating the effectiveness of contextual SoftSRV embedding. Further, comparing the performance of the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{SS}_{\mathrm{MP2}}$ methods, we observe that the more expressive parameterization of $\mathrm{SS}_{\mathrm{MC}}$ is generally beneficial. Consequently, we advocate for the use of the $\mathrm{SS}_{\mathrm{MC}}$ method in preference to other parameterized families. We focus subsequent sections on comparisons with $\mathrm{SS}_{\mathrm{MC}}$.
3.5 Comparing to Natural Language Templates
Figure 3: Gemma 2 (2B) fine-tuning curves for the synthetically generated datasets as well as the non-synthetic training set.
Here, we compare our best SoftSRV parameterization, $\mathrm{SS}_{\mathrm{MC}}$, to the two natural language prompt template methods described in Section 3.2. We first report the performance of Gemma 2B fine-tuned on the synthetic datasets generated by $\mathrm{PT}$, $\mathrm{PT}_{\mathrm{SR}}$, and $\mathrm{SS}_{\mathrm{MC}}$. Figures 3 plot the eval metrics for each dataset as a function of the number of fine-tuning steps times batch size normalized by training set size (essentially the training epoch modulo an additional constant factor due to sequence packing). We find that the model fine-tuned on $\mathrm{SS}_{\mathrm{MC}}$-generated data outperformed the models fine-tuned on data generated by the two prompt template methods.
Comparing the prompt template methods, the $\mathrm{PT}$ method outperforms the $\mathrm{PT}_{\mathrm{SR}}$ method on the GSM8K benchmark. For the MBPP benchmark, $\mathrm{PT}$ initially attains a similar performance as that of $\mathrm{PT}_{\mathrm{SR}}$, but both methods start to degrade as a function of fine-tuning steps. This may be due to a lack of diversity in the generated text, given the small set of 384 MBPP training examples used to seed the prompt templates. We do not report the results of the $\mathrm{PT}_{\mathrm{SR}}$ method for BoolQ as it appears to struggle to produce reasonable outputs. The repeated self-critiques of $\mathrm{PT}_{\mathrm{SR}}$ on the lengthy input passages seems to lead it astray from the original intention of the task, causing it to produce questions asking for open-ended discussion of a passage rather than targeted true/false questions.
We also emphasize that the training datasets from our benchmark domains contain a relatively small number of examples (see column $N_{r}$ of Table 1) and both prompt template methods and SoftSRV use this training data to generate tens of thousands of synthetic examples (see Step 2 in Section 3.3 for exact number). Yet, as shown by the results in this section, SoftSRV is more successful at leveraging this limited data. Consequently, we argue that SoftSRV is especially well suited to the data scarce setting.
Finally, as an additional comparison, we test $\mathrm{SS}_{\mathrm{MC}}$ generated data against the expertly curated GSM8K-specific synthetic dataset of Liu et al. (2023) (see detailed discussion in Section A.5). Even though differences in the synthetic generation setup introduce several confounding factors, it is nonetheless encouraging to see that the $\mathrm{SS}_{\mathrm{MC}}$-generated data is comparable to this highly curated synthetic dataset.
3.5.1 Out-of-Domain Benchmarks
Table 2: Out-of-domain task performance of Gemma 2B models fine-tuned on $\mathrm{PT}$-generated data versus on $\mathrm{SS}_{\mathrm{MC}}$-generated data, evaluated at the 2K checkpoint.
task $\mathrm{PT}$ $\mathrm{SS}_{\mathrm{MC}}$
Transcoder 0.141 0.369
MAWPS 0.583 0.547
TriviaQA 0.163 0.432
The primary focus of this work has been on generating task-specific synthetic data and, thus far, we have evaluated fine-tuned models on held-out test sets from the task-specific distribution. In this section, we measure the performance of the fine-tuned models on different, but thematically related, benchmarks to understand the generalizability of each data generation method to out-of-domain tasks. Here, we consider the following additional benchmarks: TriviaQA (Joshi et al., 2017) with a 5-shot prompt (accuracy) for reading comprehension, Transcoder (Sun et al., 2023) with a 3-shot prompt (pass@1) for coding, and MAWPS (Koncel-Kedziorski et al., 2016) with 0-shot (accuracy) for math. Each of these metrics is tested on its corresponding model, e.g. for reasoning, we evaluate on TriviaQA the model fine-tuned on the BoolQ-based synthetic datasets, etc. Table 2 shows that the $\mathrm{SS}_{\mathrm{MC}}$ method outperforms the $\mathrm{PT}$ template method on TriviaQA and Transcoder while exhibiting a closer performance on MAWPS.
Considering both in-domain results of Figures 3 and the aforementioned out-domain benchmarks, we find that, on the whole, the $\mathrm{SS}_{\mathrm{MC}}$ method exhibits superior performance compared to the $\mathrm{PT}$ template method.
3.6 Comparing to the Original Training Data
Having identified $\mathrm{SS}_{\mathrm{MC}}$ as the most promising synthetic data generation method from those we considered, we now proceed to evaluate its performance relative to the original (non-synthetic) training data from the benchmarks. Figures 3 show that the model fine-tuned on the $\mathrm{SS}_{\mathrm{MC}}$ generated data outperforms the model fine-tuned on the non-synthetic train data for MBPP and GSM8K, indicating that, given enough of it, synthetic data can outperform even non-synthetic data. However, the same observation does not hold for BoolQ. The training set curve on BoolQ admits high variance, but it attains a higher accuracy overall. As discussed in Appendix A.1, we expect generating questions for the BoolQ dataset to be more difficult both due to the broad range of question topics and also due to the relatively long context-length needed to form a good question.
3.6.1 Supplementing a post-training dataset
In the previous evaluations, we considered a setting where we fine-tune using either the synthetically generated data or the original training data. Here, we investigate augmenting an existing general SFT dataset to evaluate the effectiveness of our datasets when combined with a mixture of data sources. Specifically, we compare the performance of a 2B Gemma 2 model fine-tuned on the Gemma 2 post-training dataset (Team et al., 2024) augmented with the synthetic $\mathrm{SS}_{\mathrm{MC}}$ data versus augmented with the original training data. For all experiments, we fixed a priori the mixture source ratio to 80% for the post-training dataset and 20% for the added set (i.e. either $\mathrm{SS}_{\mathrm{MC}}$ data or training data). All other parameters remain consistent with previous sections except here the models are trained for 10K steps as the amount of data is greater.
Table 3 shows that on MBPP and GSM8K, the model fine-tuned on the mixture of Gemma 2 post-training data with $\mathrm{SS}_{\mathrm{MC}}$ generated data outperforms the model trained on the mixture of the post-training with the train data. On BoolQ, similar to our previous findings (see first paragraph of Section 3.6), the $\mathrm{SS}_{\mathrm{MC}}$-based model underperforms compared to the model using the train data.
Table 3: Downstream task performance of Gemma 2B models fine-tuned on a mixture of Gemma 2 post-training (G2) with $\mathrm{SS}_{\mathrm{MC}}$-generated data versus with the train data, evaluated at the 10K checkpoint.
task G2 + Train G2 + $\mathrm{SS}_{\mathrm{MC}}$
MBPP 0.320 0.356
GSM8K 0.423 0.539
BoolQ 0.867 0.840
3.7 Data Scaling
Next, we analyze the effects of varying the number of synthetic examples generated when fine-tuning only on synthetic examples. Specifically, we increase the number of generated examples from 20K, 50K, to 100K, testing both the $\mathrm{SS}_{\mathrm{MC}}$ and $\mathrm{PT}$ approaches for the BoolQ benchmark, given this benchmark appears the most challenging in terms of generating effective synthetic questions. Figure 4 shows that the performance of the model fine-tuned on $\mathrm{PT}$ stagnates between 20K and 50K synthetic examples, and only shows improvement when the number of generated examples reaches 100K. In contrast, the model fine-tuned on $\mathrm{SS}_{\mathrm{MC}}$ steadily improves as the generated dataset size increases. Going from 20K to 100K examples, $\mathrm{SS}_{\mathrm{MC}}$ performance increases at a 1.8 times faster rate relative to $\mathrm{PT}$ with respect to their max values. In particular, the $\mathrm{PT}$ method with 100K generated examples admits a comparable performance to that of the $\mathrm{SS}_{\mathrm{MC}}$ method with 50K generated examples.
Figure 4: We compare the BoolQ performance of Gemma 2B fine-tuned on data generated by $\mathrm{PT}$ and $\mathrm{SS}_{\mathrm{NP}}$ as the number of generated examples increases.
3.8 Distribution Matching Analysis
In addition to our evaluations of the fine-tuned model’s downstream performance, we investigate the capacity of SoftSRV and the prompt template baselines to generate data samples that can match a target distribution of text. To assess the proximity between the generated and target distributions, we compute mauve scores as in Pillutla et al. (2021), which can be viewed as a scalar summary of the divergence between the textual output of a generative model and a reference distribution. The mauve score is able to simultaneously measure both the model’s ability to avoid generating text outside the support of the target distribution (Type I error) and the ability to generate text with a large coverage of the target distribution support (Type II error). The method essentially computes a quantized distribution for the generated and target distribution and measures their KL divergence, producing a normalized score between 0 and 1, where 1 indicates the two distributions are maximally similar (for further details see Appendix A.7).
Table 4: The mauve scores of synthetic datasets computed with respect to the non-synthetic test fold of each dataset.
MBPP GSM8K BoolQ
$\mathrm{PT}$ 0.463 0.914 0.784
$\mathrm{PT}_{\mathrm{SR}}$ 0.397 0.865
$\mathrm{SS}_{\mathrm{MC}}$ 0.477 0.991 0.995
Train 0.963 0.998 0.999
In Table 4, we report the mauve scores for the template prompt methods and for $\mathrm{SS}_{\mathrm{MC}}$ on the MBPP, GSM8K and BoolQ domains, measuring the distance to questions in the test fold (distance to question in the train fold are shown in Appendix A.7). In all cases, we find that $\mathrm{SS}_{\mathrm{MC}}$ can synthesize text that is closer to a sample from the target distribution than the prompt template approaches. Notably, the $\mathrm{SS}_{\mathrm{MC}}$ method achieves a high mauve score that is even close to that of the train set on GSM8K and BoolQ, where it appears the additional flexibility afforded by its parameterization allows a very high-fidelity match to the target distribution. While the prompt template-based datasets generally have lower mauve scores than $\mathrm{SS}_{\mathrm{MC}}$, the $\mathrm{PT}$ variant is able to achieve a relatively high score on GSM8K. In the case of the MBPP dataset, we measure very low similarity scores for $\mathrm{PT}_{\mathrm{SR}}$; we conjecture this may be due to the several rounds of rewriting, which results in straying further from the original seed question.
Even though the mauve score is not a direct indicator of downstream fine-tuning performance (see our experiments in Section 3.5 for performance comparisons), it is an additional signal to provide further support for SoftSRV as a high-fidelity approach for text generation across domains.
4 Related Work
As mentioned in the introduction, there is a significant recent body of work demonstrating the effectiveness of using a LLM to generate synthetic training data for a smaller model.
In terms for generating pre-training data, the collection of “Textbooks Are All You Need” white-papers outlines the process of training the Phi series of small LMs, using carefully curated prompt templates and seed data sources, and shows the large boost in quality that synthetic data can provide (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024). The Cosmopedia project (Ben Allal et al., 2024) conducts a similar study, while open-sourcing the prompts and generated data. Li et al. (2024a) construct a pretraining set almost “from scratch” using a diverse set of prompt templates by using taxonomies of fields/sub-fields/disciplines within an area of expertise. Apart from focusing on pre-training rather than fine-tuning, these works require a non-trivial amount of human effort for building and/or curating prompt templates for the generating LLM, which our effort seeks to minimize.
Finally, Lee et al. (2024) recently proposed an adaptive procedure where a LLM is used to generate targeted fine-tuning data for a small model based on examples that the small model has made mistakes on. The LLM is prompted to rewrite variants of these questions using specialized per-domain prompt templates. Extending SoftSRV to an adaptive setting is a potentially promising future line of work.
5 Conclusion
In this work, we established the effectiveness of the SoftSRV framework for generating targeted synthetic fine-tuning data. We deploy the same SoftSRV pipeline across math, coding, and reasoning tasks, finding in each case that SoftSRV generates fine-tuning data that provides strong downstream performance with no manual prompt engineering or per-domain specialization needed. A potentially fruitful direction for future work is to adaptively select or generate context vectors $\mathbf{z}$, in order to generate even more impactful synthetic data.
Appendix A Empirical Evaluation Additional Details
A.1 Benchmark Task Discussion
We consider the following tasks:
@@ A.1 Benchmark Task Discussion
Reasoning – BoolQ (Clark et al., 2019). For general reasoning, we consider the binary question answering dataset from the BoolQ benchmark. The questions arise organically from anonymized Google search queries which can be answered as either ‘true’ or ‘false’. Each question and answer is paired with a passage (average length of 108 tokens) that is extracted from a relevant Wikipedia page. This is evaluated as a scoring/classification task and accuracy is reported.
@@
The MBBP task requires basic Python programming knowledge to solve, but the questions are generally short, follow a similar pattern and are concerned with a relatively narrow set of themes. The GSM8K task requires basic math and language comprehension skills, but generating the problem is arguably even harder than solving it. It requires generating a premise containing several numerical quantities and then a question that can be answered using the provided information in a non-trivial fashion. Nonetheless the premises are somewhat formulaic and thematically similar.
The BoolQ reasoning task, which requires general reading comprehension to solve, is perhaps the most difficult task to generate synthetic data for. Generating a problem requires writing a long (relative to MBPP and GSM8K) passage on an arbitrary topic that contains a collection of facts, but that does not necessarily stick to any formula or theme, and then generate a true/false question that can be answered directly by the passage. As we shall see in the empirical evaluation, the level of difficulty in generating a high-quality question can impact the relative quality and value of generated synthetic data.
A.2 Details on “Step 2. Generate Questions”
@@
In the case of SoftSRV, we cycle through questions found in the training fold of the benchmark dataset, $x_{i}$, and produce a new synthetic questions $x_{i}^{\prime}\sim H(\mathbf{P}_{\theta}(E(x_{i})))$ via temperature sampling (with default temp=1). For the $\mathrm{SS}_{\mathrm{NP}}$ variant, no context vector and, thus, no training examples are used during generation. For all SoftSRV methods, no natural language prompt template of any kind is used.